OpenStudy (anonymous):

Choose the equation of the line passing through the point (-2, 6) and parallel to y = x - 8. y = x + 7 y = x + 5 y = 2x + 7 y = 2x + 5 help please ?

OpenStudy (anonymous):

slope wiill be 1. so by formula, (y-y1)=m(x-x1) (y-6)=1.(x-(-2)) nowsolve it

OpenStudy (anonymous):

have u rwritten the ques and options correctly?

OpenStudy (anonymous):

yes is it the 3rd one ?

OpenStudy (anonymous):

my ans is comin y=x+8

OpenStudy (anonymous):

thats weird thats not one of the options maybe were doing something wrong ???

OpenStudy (anonymous):

given eqn is y=x-8 so slope will me 1. and,slope of reuired line will be also 1 bcon it is paralled. and the given point is (-2,6) now jst we have to put it in the formula....dios is the process

OpenStudy (anonymous):

oh lol so how would we put it in the formula?

OpenStudy (anonymous):

(y-y1)=slope(x-x1) x1 and yi are the given point`

OpenStudy (anonymous):

so than wouldn't it be the last one ?

OpenStudy (anonymous):

@madrockz

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

well than which one would it be ?

OpenStudy (anonymous):

the 2nd on ?

OpenStudy (anonymous):

omg i just noticed something I'm such an idiot lmao the correct question and answer are Choose the equation of the line passing through the point (-2, 6) and parallel to y =1/2 x - 8. y = 1/2x + 7 y = 1/2x + 5 y = 2x + 7 y = 2x + 5 @madrockz

OpenStudy (anonymous):

see i told u to check it once ur ques.my procwess is right.

OpenStudy (anonymous):

lol i know I'm sorry

OpenStudy (anonymous):

so now which one do you think it is ?

OpenStudy (anonymous):

its ok...:-)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

i told u the process u try it i will guide u

OpenStudy (anonymous):

ok now tell me wat will be the slope of required line

OpenStudy (anonymous):

so nothing changed in the process now that the numbers changes ?

OpenStudy (anonymous):

um idk 1 ??

OpenStudy (anonymous):

sorry I'm not good at algebra

OpenStudy (anonymous):

the coefficient of x is always the slope. y=mx+c where m is the slope of line.do u know dis?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

ok so look at the equation and tell me the slope now.

OpenStudy (anonymous):

wait is the slope a number or a letter ? sorry I'm not the best at this

OpenStudy (anonymous):

no.is the slope...the no. astuck wid x is the slpe now tell me

OpenStudy (anonymous):

1/2 is that it ?

OpenStudy (anonymous):

yes very good.it is the slpe.

OpenStudy (anonymous):

yay lol so what now ?

OpenStudy (anonymous):

now the equation to find out when a point and a slopw is give....is represented by (y-y1)=m(x-x1) where m is the slpoe and wt is the given poin t tell me

OpenStudy (anonymous):

is the given point 8 ??

OpenStudy (anonymous):

no..the given point is (-2,6).isnt it?have a look at the ques

OpenStudy (anonymous):

yes it is my bad lol

OpenStudy (anonymous):

-2 is x1 and 6 is y1 now put it in theequation which i told u

OpenStudy (anonymous):

(Y-6y)= 1/2(x-(-2x)) is that good?

OpenStudy (anonymous):

it is a good....bt u have to omit x and y u only have to write the no.s ok.now remove the x and

OpenStudy (anonymous):

ok let me tell u (y-6)=1/2(x-(-2)) ok??

OpenStudy (anonymous):

6y=1/2(x-(-2))

OpenStudy (anonymous):

oh yess

OpenStudy (anonymous):

u got it?sure?

OpenStudy (anonymous):

bess thank you so much for everything

OpenStudy (anonymous):

ok then now we have to simplify it

OpenStudy (anonymous):

can u try it once?

OpenStudy (anonymous):

so would the answer be the 1st answer or no ?

OpenStudy (anonymous):

u have to simplify it then u will get the ans

OpenStudy (anonymous):

2(y-6)=1.(x+2) do u agree? dis step?

OpenStudy (anonymous):

see i will help u..u jst have to have the will to learn

OpenStudy (anonymous):

yes i agree so we solve it from there right ?

OpenStudy (anonymous):

yes multiply the bracket wid 2...can u do it nd show me?

OpenStudy (anonymous):

i got 12 when i multiplied the bracket with 2 is that right

OpenStudy (anonymous):

so would it be (12y) =1.(x+2 ?

OpenStudy (anonymous):

when u multiply u will get 2y-12 ok

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

do we do the something for the 1.(x+2)

OpenStudy (anonymous):

see when u multiply 2 wid y u get 2y and when u multiply 2 wid 6 u get 12

OpenStudy (anonymous):

yess i see

OpenStudy (anonymous):

yes the eqn will v now, 2y-12=x+2 2y-x=2+12

OpenStudy (anonymous):

now can tu tell me wt will be the ans

OpenStudy (anonymous):

first solve it completely add 2 and 12

OpenStudy (anonymous):

show me wt u got finally

OpenStudy (anonymous):

okay so 2y - x = 14

OpenStudy (anonymous):

yes very good

OpenStudy (anonymous):

now we have to bring dis eqn in form of option

OpenStudy (anonymous):

How would we do that ?

OpenStudy (anonymous):

take the x to the left side.can u do it?

OpenStudy (anonymous):

so like x=2y +14 or in a different way?

OpenStudy (anonymous):

no...not in dis way see eqn was 2y-x=14 then 2y=x+14 y=x/2+7 now chck it ans

OpenStudy (anonymous):

so its the first one ?

OpenStudy (anonymous):

yes first one

OpenStudy (anonymous):

thanks you

OpenStudy (anonymous):

welcome!!!